# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

### v1

@@ Abstract
Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models. However, they maintain additional moving average states throughout training, which results in memory requirements several times greater than the model. This overhead imposes constraints on scalability and computational efficiency. On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency, their capability in LLM training is limited (Zhao et al., 2024b).
To address this dilemma, we show that pre-processing SGD is sufficient to reach Adam-level performance on LLMs. Specifically, we propose to preprocess the instantaneous stochastic gradients with two simple operators: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ stabilizes gradient distributions, and $\mathtt{GradWhitening}$ counteracts the local curvature of the loss landscape, respectively. This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any accumulative state variables. Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam. In language modeling tasks, SWAN demonstrates the same or even a substantial improvement over Adam. Specifically, when pre-training the LLaMa model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity in less than half tokens seen.
@@ 1 Introduction
Adaptive optimizers, such as Adam and its variants (Kingma & Ba, 2015; Loshchilov & Hutter, 2019; Shazeer & Stern, 2018; Pagliardini et al., 2024; Liu et al., 2023; Zhao et al., 2024a), have been central to the success of training large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; Bi et al., 2024; Bai et al., 2023; Zhang et al., 2022). However, most adaptive optimizers for LLMs are stateful, meaning they require tracking and maintaining internal states. While achieving remarkable empirical success, these states introduce significant memory overhead. For instance, Adam (Kingma & Ba, 2015)–the de facto optimizer for LLM training–involves the tracking of exponential moving averages (EMAs), effectively doubling memory requirements. AdEMAMix (Pagliardini et al., 2024) – an extension of Adam that achieves significant convergence speed boost – requires storing even more states, tripling the memory requirements. This overhead can be significant especially in distributed settings, where the optimizer states could consume a significant amount of the GPU memory (Dubey et al., 2024; Korthikanti et al., 2023). On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency (i.e., they are stateless), their capability to train LLMs is limited (Zhao et al., 2024b), and the performance gap between SGD and Adam is significant (Zhang et al., 2020; Kunstner et al., 2023; 2024).
To address the memory-performance dilemma for LLMs, this work questions the requirement of accumulative state variables during training to achieve good performance in LLMs. We propose to simply preprocess the instantaneous SGD gradient in a history-independent manner, such that it still preserves certain desirable properties of adaptive optimizers. This result in SWAN (SGD with Whitening And Normalization), a novel SGD-like optimizer that eliminates all internal optimizer states and empirically surpasses the performance Adam on several LLM pre-training scenarios.
@@
Our proposed optimizer consists of two stateless operators applied to the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ aims to stabilize the stochasticity of gradient distributions during training, while $\mathtt{GradWhitening}$ seeks to neutralize the local geometry of the loss landscape. Both operators enables SWAN to rely solely on the statistics of the current gradient matrices. This SGD preprocessing approach eliminates all internal optimizer states, and therefore reproduces the memory consumption of SGD, reaching a new state-of-the-art in terms of optimizer memory efficiency for LLMs ($\approx 50\%$ reduction on end-to-end memory footprint vs Adam, and $\approx 100\%$ reduction in optimizer states. See Figure 1). In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA models on the C4 dataset with multiple model sizes show consistently better performance than Adam and other low-rank memory-efficient optimizers. Remarkably, at the 350M and 1.3B scale, our method achieves 2X faster convergence in terms of tokens seen compared to Adam, marking SWAN the first known stateless optimizer converges as fast/faster than Adam in training large language models.
@@
Figure 1: SWAN performance preview on LLM pretraining. (a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves $>$ 2X speed-up vs Adam in terms of tokens seen. (c): memory footprint. Unlike Zhao et al. (2024a), we directly measure end-to-end memory under full-model training, with batch size = 1 sequence. SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states). (d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam. SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
Our contributions are as follows:
[…]
    1. 1.
      Robustness to ill-conditioned problems. On LLM pretraining, SWAN converges quickly even without learning rate warm-ups (Section 5.3, Section 5.6), a property not observed in Adam.
    2. 2.
    3. 3.
      Computational efficiency: consistently outperforms Adam and reaches up to 2$\times$ acceleration and effective throughput.
  • Theoretical analysis of SWAN. We show that (1) $\mathtt{GradNorm}$ can be derived as a strategy to stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 4.3). Additionally, we highlight that $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 3, Section B.3), offering theoretical advantages over SGD and Adam (Proposition 3).
  • Fast convergence for LLM pretraining tasks. Through experiments (Section 5) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also outperforms Adam and other memory-efficient optimizers. Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
@@ 2 Preliminaries
One of the most used adaptive optimizers is the Adam optimizer (Kingma & Ba, 2015). Adam is an example of stateful optimizers, which involves tracking and maintaining internal states. It combines the advantages of two earlier methods: AdaGrad (Duchi et al., 2011), which adapts learning rates based on the historical gradient magnitude, and RMSProp (Tieleman, 2012), which mitigates the aggressive decrease in learning rates by using a decaying average of squared gradients.
@@
where ${\bm{m}}^{(t)}$ and ${\bm{\nu}}^{(t)}$ are EMAs of the first and second moments of the gradients; and $\eta$ is a global step size. Essentially, Adam estimates the signal-to-noise (SNR) ratios and use it to adjust learning rates element-wise. Tracking and storing these two EMA estimates triple the total memory consumption required to train a LLM model. For example, for LLaMA 405B model, storing model weights requires 810 GB of memory, while the Adam optimizer states requires an additional 1.6TB of memory.
@@
There is a rich literature on understanding adaptive methods’ inner workings and unreasonable effectiveness. Notably, the key desired properties of Adam can be categorized as gradient smoothing, gradient invariance and gradient whitening. (1) Gradient smoothing: the EMA operations in Adam naturally reduce the influence of mini-batch noise Cutkosky & Mehta (2020); Crawshaw et al. (2022); (2) Gradient invariance: recent work suggest the performance gap between SGD and Adam might lie in Adam’s sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023). For example, Adam is robust to the rescaling of gradient diagonals (Kingma & Ba, 2015); and is invariant to any sign-preserving scalings (under $\beta_{1}=\beta_{2}=0$)(Bernstein et al., 2018). And (3) Gradient whitening : it is well-known that the second moment EMA of Adam approximates the diagonal of Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024), and biases the optimization trajectories towards well-conditioned regions (Jiang et al., 2024). For a more comprehensive discussion of these properties, please refer to appendix A.
Adam as SGD preprocessing, and the potential unnecessity of EMAs
Adam can be viewed as history-dependent preprocessing of the gradients ($\{{\mathbf{G}}^{(0)},{\mathbf{G}}^{(1)},...,{\mathbf{G}}^{(t-1)},{\mathbf{G}}^{(t)}\}\rightarrow\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$) to achieve the desired properties above. A key observation is, all of the above properties are achieved through the element-wise operations. That is, each element of the gradient matrix is pre-processed and re-scaled independently. This ignores the interactions and structures among different variables, and we hypothesize this is the reason why extra history information is need to compensate this gap, leading to the requirement of EMA states. Thus, we believe that designing stateless adaptive optimizers is possible, if we can achieve similar properties by designing new matrix-level operations that are preprocesses the instantaneous gradients of SGD (${\mathbf{G}}^{(t)}\rightarrow\tilde{{\mathbf{G}}}^{(t)}$).
@@ 3 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
Figure 2: Illustration of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ operators. In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise. In $\mathtt{GradWhitening}$ operator, we treat each column (or row) of the gradient matrix ${\mathbf{G}}$ as a separate data sample. Then, the $\mathtt{GradWhitening}$ operator can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is an identity matrix. Note that the same approach can be applied row-wise as well.
@@
  Input: weight matrix ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ with $m\leq n$. Step size $\eta$. Number of $\mathtt{GradWhitening}$ iteration $k$.
@@
  Input: Matrix ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of Newton Schultz iteration $k$, step size $\beta$
@@
  Input: Matrix ${\mathbf{G}}^{m\times n}$
  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\sum_{j=1}^{n}{\mathbf{G}}_{:,j})^{2}}\mathbf{1}_{n}^{\top}}$
Figure 3: SWAN Optimizer. Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximant. Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
As discussed in Section 2, we believe the key to designing stateless, adaptive, and effective optimizers lies in designing new matrix-level operations that exploit rich information in the gradient matrix. To this end, we propose a new stateless optimizer (Algorithm 1), SWAN (SGD with Whitening And Normalization). SWAN preprocesses SGD based on two simple operators, $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$, both are stateless, matrix-level operators. When applied in tandem, they achieve similar desirable properties of adaptive optimizers, without the need to store historical gradient moments. In this section we describe the update rules of SWAN.

### v2

# SWAN: SGD with Normalization and Whitening Enables Stateless LLM Training

@@ Abstract
Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models. However, they often require maintaining optimizer states throughout training, which can result in memory requirements several times greater than the model footprint. This overhead imposes constraints on scalability and computational efficiency. Stochastic Gradient Descent (SGD), in contrast, is a stateless optimizer, as it does not track state variables during training. Consequently, it achieves optimal memory efficiency. However, its capability in LLM training is limited (Zhao et al., 2024b). In this work, we show that pre-processing SGD in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while drastically reduces the memory cost. Specifically, we propose to pre-process the instantaneous stochastic gradients using normalization and whitening. We show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape. This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any optimizer states. Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam. In language modeling tasks, SWAN demonstrates comparable or even better performance than Adam: when pre-training the LLaMA model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity using half as many tokens.
@@ 1 Introduction
Adaptive optimizers, such as Adam and its variants (Kingma & Ba, 2015; Loshchilov & Hutter, 2019; Shazeer & Stern, 2018; Pagliardini et al., 2024; Liu et al., 2023; Zhao et al., 2024a), have been central to the success of training large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023b; Dubey et al., 2024; Bi et al., 2024; Bai et al., 2023; Zhang et al., 2022). However, most adaptive optimizers for LLMs are stateful, meaning they require tracking and maintaining internal states. While achieving remarkable empirical success, these states introduce significant memory overhead. For instance, Adam (Kingma & Ba, 2015) – the de facto optimizer for LLM training – involves the tracking of exponential moving averages (EMAs), effectively doubling memory requirements. AdEMAMix (Pagliardini et al., 2024), an extension of Adam that achieves significant convergence speed boost, requires storing even more states, tripling the memory requirements. This overhead can be significant especially in distributed settings, where the optimizer states could consume a significant amount of the GPU memory (Dubey et al., 2024; Korthikanti et al., 2023). On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency (i.e., it is stateless), their capability to train LLMs is limited (Zhao et al., 2024b; Zhang et al., 2020; Kunstner et al., 2023; 2024). Therefore, a natural question arises:
Can LLMs be trained efficiently using stateless optimizers?
There is a growing body of research that has contributed to answering this question positively by developing novel optimizers that reduce the memory requirements associated with tracking internal state variables during the training of LLMs while achieving similar or even speedup boost performance compared to Adam. For instance, some methods rely solely on tracking the first moment of gradients (Xu et al., 2024a; Jordan et al., 2024), while others introduce an additional one-dimensional tracking variable on top of first moments (Zhang et al., 2024b; Zhao et al., 2024c). Alternatively, approaches focusing exclusively on pre-conditioner tracking have also been proposed (Pooladzandi & Li, 2024; Li, 2017). Another line of work focuses on using low-rank approximations to store the first and second moments, thereby reducing the memory cost associated with tracking optimizer states (Lialin et al., 2023; Hao et al., 2024; Zhao et al., 2024a). More recently, Zhu et al. (2024) introduced an almost stateless optimizer for efficient LLM training, which proposes to only track rank-1 state variables during training.
In this work, we address this question by proposing to simply pre-process the instantaneous stochastic gradient in a stateless manner. The result is SWAN (SGD with Whitening And Normalization), a novel stochastic optimizer that eliminates all internal optimizer states and empirically achieves comparable or even better performance compared to Adam on several LLM pre-training tasks.
@@
Our optimizer consists of combining two well-known operators to pre-process the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ applies a row-wise standardization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix. We show that these operators aims at stabilizing the stochasticity of gradient distributions during training, and neutralizing the local geometry of the loss landscape, respectively. When applied together, both operators enables SWAN to rely solely on the statistics of the current gradient matrices. This approach eliminates the need to track state variables, thereby matching the memory footprint of SGD. In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA (Touvron et al., 2023a) models on the C4 dataset with multiple model sizes show consistently the same or better performance than Adam and other low-rank optimizers. Remarkably, at the 350M and 1.3B scale, our method achieves up to 2X faster convergence in terms of tokens seen compared to Adam. Our contributions are summarized below:
[…]
    1. 1.
      Robustness to ill-conditioned problems: on LLM pre-training, SWAN consistently and efficiently converges without learning rate warm-up (Section 6.1, Section 6.4).
    2. 2.
    3. 3.
      Computational efficiency: consistently achieves the same or better performance when compared to Adam across several scales and can significantly improve token-efficiency during training up to 2$\times$ at the 1B scale for the LLaMA model.
  • Theoretical analysis of SWAN. We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3). Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section B.1).
  • Fast convergence for LLM pretraining. Through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam. Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
@@
Figure 1: SWAN performance preview on LLM pretraining. (a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves $>$ 2X speed-up vs Adam in terms of tokens seen. (c): memory footprint. We directly measure end-to-end memory under full-model training, with batch size = 1 sequence. This is different to the layer-wise training setting Zhao et al. (2024a). SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states). (d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the token efficiency of optimizer relative to Adam. SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
2 Related works
Towards Stateless Optimizers for LLM Training.
Adaptive optimizers generally rely on tracking internal state variables to perform weight updates, which can substantially increase memory consumption when training large models. Several recent works have successfully managed to reduce the memory requirements associated with storing additional state variables for training LLMs. Muon (Jordan et al., 2024; Bernstein & Newhouse, 2024b), a newly proposed optimizer, has demonstrated strong acceleration and memory saving for LLM training by simplifying shampoo-like optimizers (Gupta et al., 2018; Anil et al., 2021; Shi et al., 2023; Wang et al., 2024; Vyas et al., 2024; Peirson et al., 2022; Lin et al., 2024) and requiring only the tracking of a first-moment estimate. SGD-Sal (Xu et al., 2024b) only stores the first moment with a learning rate estimated at the beginning of training. Sign-based methods such as (Chen et al., 2023) have also demonstrated success on training transformer-based models by only tracking first moments. There are also several works that aim to enhance the memory efficiency of Adam by reducing the memory cost associated with second moments. Adam-mini (Zhang et al., 2024b) significantly reduces memory usage by storing only scalar values for each parameter block, while Adalayer (Zhao et al., 2024c) retains only the scalar average of the second moment for each layer. Alternatively, PSGD (Pooladzandi & Li, 2024; Li, 2017) focuses on exclusively tracking a pre-conditioner, eliminating the need to track a first moment estimate. Finally,  (Zhu et al., 2024) proposes an nearly stateless optimizer to train LLMs, which only requires tracking rank-1 first and second moments. However, all the aforementioned optimizers still require the storage of state variables. In contrast, SWAN completely eliminates the need to store internal states for both the first and second moments by employing a combination of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ steps, which is discussed next.
Pre-processing Gradients.
Gradient pre-processing is a common technique used to enhance performance of optimizers. Various pre-processing procedures have been proposed in the literature, such as signed gradient (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023; Kunstner et al., 2023), gradient clipping (Zhang et al., 2020), normalization (Zhang et al., 2020; You et al., 2019), and whitening (Yang & Laaksonen, 2008; Kingma & Ba, 2015; Hwang, 2024; Jordan et al., 2024; Bernstein & Newhouse, 2024c; a; Carlson et al., 2015). In this work, we particularly focus on normalization and whitening. We apply normalization row-wise on gradient matrices, similar to (You et al., 2019); together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022). Our key result is that composing normalization and whitening on stochastic gradients is sufficient to enable the efficient training of LLMs in a completely stateless manner. Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients instead of Adam states. Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses normalization. We show empirically that removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4).
Low-rank methods.
Low-rank optimization techniques have been explored in the context of large language model (LLM) training as a means to reduce memory consumption. These methods focus on applying low-rank approximations to model weights, gradients, and/or optimizer state variables. A seminal work in this domain is LORA (Hu et al., 2021) to fine-tune pre-trained models using additional low-rank weight matrices at each layer, thereby significantly reducing memory usage to update the weights. More recently, methods such as ReLoRA (Lialin et al., 2023), FLORA (Hao et al., 2024), and Galore (Zhao et al., 2024a) have advanced low-rank optimization techniques for memory-efficient LLM pre-training. These approaches leverage low-rank gradient projections to enable full-rank learning, thereby achieving memory savings without compromising model capacity. Notably, they have achieved substantial reductions in the memory consumption of optimizer states, with only minimal impact on model performance. While these approaches effectively reduce the memory footprint of LLM training, they still necessitate storing internal states, resulting in higher memory consumption compared to SWAN.
@@ 3 Preliminaries
Adam (Kingma & Ba, 2015) is the current standard choice for adaptive optimizers across a multitude of machine learning training tasks, including LLM pre-training. Adam is an example of a stateful optimizer, which means it accumulates and stores internal states throughout training. It combines the advantages of two earlier methods: AdaGrad (Duchi et al., 2011), which adapts learning rates based on the historical gradients’ magnitudes, and RMSProp (Tieleman, 2012), which mitigates the aggressive decrease in learning rates by using a decaying average of squared gradients.
@@
where ${\bm{m}}^{(t)}$ and ${\bm{\nu}}^{(t)}$ are EMAs of the first and second moments of the gradients; and $\eta$ is a global step size. Intuitively, Adam estimates the signal-to-noise (SNR) ratios and use it to adjust learning rates element-wise. Tracking and storing these two EMA estimates triple the total memory consumption required to train a LLM model. For example, for LLaMA 405B model, storing model weights requires 810 GB of memory, while the Adam optimizer states requires an additional 1.6TB of memory.
@@
There is a rich literature on understanding adaptive methods’ inner workings and unreasonable effectiveness. Notably, the key desired properties of Adam can be categorized as gradient whitening, gradient smoothing, and gradient invariance.
  • Gradient whitening: it is known that the inverse second moment $\frac{1}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$ of Adam performs gradient whitening by approximating the square root inverse of the diagonal of Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024). This step biases the optimization trajectories towards well-conditioned regions (Jiang et al., 2024) and provides a better approximation to the geodesic flow when compared with the natural gradient update (Yang & Laaksonen, 2008).
  • Gradient smoothing: the EMA operations in Adam naturally reduce the influence of mini-batch noise (Cutkosky & Mehta, 2020; Crawshaw et al., 2022);
  • Gradient invariance: recent work suggest the performance gap between SGD and Adam might lie in Adam’s sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023). For example, Adam is robust to the rescaling of gradient diagonals (Kingma & Ba, 2015); and is invariant to any sign-preserving scalings (under $\beta_{1}=\beta_{2}=0$(Bernstein et al., 2018).
For a more comprehensive discussion of these properties, please refer to appendix A.
Adam as SGD pre-processing.
Adam can be viewed as history-dependent pre-processing of the gradients ($\{{\mathbf{G}}^{(0)},{\mathbf{G}}^{(1)},...,{\mathbf{G}}^{(t)}\}\rightarrow\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$) to achieve the desired properties described above. A key observation is that all of these properties are achieved through element-wise operations, where each element of the gradient matrix is independently pre-processed and re-scaled. This approach does not take into account the interactions and structures between different variables, and we hypothesize that this is the reason why additional history information is necessary to bridge this gap, ultimately leading to the requirement for EMA states. Thus, we believe that designing stateless adaptive optimizers is possible if we can achieve similar properties by applying matrix-level operations that pre-process the instantaneous stochastic gradients of SGD (${\mathbf{G}}^{(t)}\rightarrow\tilde{{\mathbf{G}}}^{(t)}$).
@@ 4 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
Figure 2: Illustration of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ operators. In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise. In $\mathtt{GradWhitening}$ operator (illustration adapted from Huang et al. (2019)), we treat each column of the gradient matrix ${\mathbf{G}}$ as a separate data sample. Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is the identity matrix.
@@
  Input: weight matrix ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ with $m\leq n$. Step size $\eta$. Number of $\mathtt{GradWhitening}$ iteration $k$ (default = 10).
@@
  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $k$. Step size $\beta$ (default = 0.8).
@@
  Input: ${\mathbf{G}}^{m\times n}$
  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\frac{1}{n}\sum_{k=1}^{n}{\mathbf{G}}_{:,k})^{2}}\mathbf{1}_{n}^{\top}}$
Figure 3: SWAN Optimizer. Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximate Song et al. (2022). Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
As discussed in Section 3, we believe the key to designing stateless, adaptive, and effective optimizers lies in the incorporation of matrix-level operations that exploit rich information contained in the gradient matrix. To this end, we compose two well-known matrix operators, namely normalization and whitening. When applied in tandem, they achieve similar desirable properties of adaptive optimizers, without the need to store historical gradient moments. The result is SWAN (SGD with Whitening And Normalization), a new stateless optimizer which we describe next.